how many pounds of pure chocolate must be added to 20 pounds of brownie dough that is 30% chocolate to produce a brownie mixture that is 50% chocolate?
Answer by stanbon(75887)

how many pounds of pure chocolate must be added to 20 pounds of brownie dough that is 30% chocolate to produce a brownie mixture that is 50% chocolate?
-----------------------------------
One variable Solution:
chocolate + chocolate = chocolate:
0.30*20 + 1.00x = 0.50(20+x)
Multiply thru by 100 to get:
30*20 + 100x = 50*20 + 50x
50x = 20*20
=================
x = 8 lbs (amt. of pure chocolate that must be added to the mixture)
============================
Cheers,
Stan H.